OpenStudy (anonymous):

What is the solution set of |x – 6| = 4?

OpenStudy (anonymous):

x= -10,-2

OpenStudy (anonymous):

2 and 10

OpenStudy (anonymous):

Erm i mean 2 and 10

OpenStudy (anonymous):

x-6=4 or x-6=-4 x=10 or x=2

OpenStudy (anonymous):

:P